हिंदी

The probability that a student will pass the final examination in both English and Hindi is 0.5 and the probability of passing neither is 0.1.

Advertisements
Advertisements

प्रश्न

The probability that a student will pass the final examination in both English and Hindi is 0.5 and the probability of passing neither is 0.1. If the probability of passing the English Examination is 0.75. What is the probability of passing the Hindi Examination?

योग
Advertisements

उत्तर

Let A and B be the events of passing English and Hindi examinations, respectively.

Accordingly, we have:

P(A and B) = 0.5

 P(not A and not B) = 0.1 [i.e. P(A' ∩ B') = 0.1]

P(A) = 0.75

Now, P(A∪B) + P(A' ∩ B') = 1

⇒ P(A∪B) = 1 - P(A' ∩ B')

                  = 1 -0.1 = 0.9

By addition theorem, we have:

P(A ∪ B) = P(A) + P(B) - P(A ∩ B)

⇒ 0.9 = 0.75 + P (B)  - 0.5

⇒ P(B) = 0.9 - 0.75 + 0.5

⇒ P(B) = 0.65

Thus, the probability of passing the Hindi examination is 0.65.

shaalaa.com
Concept of Probability - Probability of 'Not', 'And' and 'Or' Events
  क्या इस प्रश्न या उत्तर में कोई त्रुटि है?
अध्याय 14: Probability - EXERCISE 14.2 [पृष्ठ ३०७]

APPEARS IN

एनसीईआरटी Mathematics [English] Class 11
अध्याय 14 Probability
EXERCISE 14.2 | Q 20. | पृष्ठ ३०७
आर.डी. शर्मा Mathematics [English] Class 11
अध्याय 33 Probability
Exercise 33.4 | Q 13 | पृष्ठ ६८

संबंधित प्रश्न

If `2/11` is the probability of an event, what is the probability of the event ‘not A’.


A letter is chosen at random from the word ‘ASSASSINATION’. Find the probability that letter is

  1. a vowel
  2. an consonant

A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16. Determine P(A or B).


In Class XI of a school 40% of the students study Mathematics and 30% study Biology. 10% of the class study both Mathematics and Biology. If a student is selected at random from the class, find the probability that he will be studying Mathematics or Biology.


The probability that a student will pass the final examination in both English and Hindi is 0.5 and the probability of passing neither is 0.1. If the probability of passing the English examination is 0.75, what is the probability of passing the Hindi examination?


A die has two faces each with number ‘1’, three faces each with number ‘2’ and one face with number ‘3’. If die is rolled once, determine

  1. P(2)
  2. P(1 or 3)
  3. P(not 3)

In a certain lottery, 10,000 tickets are sold and ten equal prizes are awarded. What is the probability of not getting a prize if you buy one ticket.


Three letters are dictated to three persons and an envelope is addressed to each of them, the letters are inserted into the envelopes at random so that each envelope contains exactly one letter. Find the probability that at least one letter is in its proper envelope.


A dice is thrown. Find the probability of getting:

 2 or 4


In a simultaneous throw of a pair of dice, find the probability of getting:

8 as the sum


In a simultaneous throw of a pair of dice, find the probability of getting  an even number on first


In a simultaneous throw of a pair of dice, find the probability of getting an even number on one and a multiple of 3 on the other


In a simultaneous throw of a pair of dice, find the probability of getting neither 9 nor 11 as the sum of the numbers on the faces


In a simultaneous throw of a pair of dice, find the probability of getting a total greater than 8.

 

In a single throw of three dice, find the probability of getting a total of 17 or 18.

 

Three coins are tossed together. Find the probability of getting exactly two heads


Three coins are tossed together. Find the probability of getting at least one head and one tail.

 

A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is white .


A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is white and odd numbered .


A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is red or even numbered.


Fill in the blank in the table:

P (A) P (B) P (A ∩ B) P(A∪ B)
\[\frac{1}{3}\] \[\frac{1}{5}\] \[\frac{1}{15}\] ......

Fill in the blank in the table:

P (A) P (B) P (A ∩ B) P(A∪ B)
0.35 .... 0.25 0.6

If A and B are two events associated with a random experiment such that
P (A ∪ B) = 0.8, P (A ∩ B) = 0.3 and P \[(\bar{A} )\]= 0.5, find P(B).

 


There are three events A, B, C one of which must and only one can happen, the odds are 8 to 3 against A, 5 to 2 against B, find the odds against C


One of the two events must happen. Given that the chance of one is two-third of the other, find the odds in favour of the other.


A card is drawn from a deck of 52 cards. Find the probability of getting an ace or a spade card.


100 students appeared for two examination, 60 passed the first, 50 passed the second and 30 passed both. Find the probability that a student selected at random has passed at least one examination.


The probability that a leap year will have 53 Fridays or 53 Saturdays is


A and B are two events such that P (A) = 0.25 and P (B) = 0.50. The probability of both happening together is 0.14. The probability of both A and B not happening is


If the probability of A to fail in an examination is \[\frac{1}{5}\]  and that of B is \[\frac{3}{10}\] . Then, the probability that either A or B fails is

 
 

A box contains  10 good articles and 6 defective articles. One item is drawn at random. The probability that it is either good or has a defect, is


A bag contains 5 black balls, 4 white balls and 3 red balls. If a ball is selected randomwise, the probability that it is black or red ball is


Two dice are thrown simultaneously. The probability of getting a pair of aces is


Three numbers are chosen from 1 to 20. The probability that they are not consecutive is


In a certain lottery 10,000 tickets are sold and ten equal prizes are awarded. What is the probability of not getting a prize if you buy two tickets.


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×